OpenStudy (anonymous):

Rachel buys a new television that is on sale for 15% off. She also gets a 5% discount on the sale price plus a $50 mail-in rebate. If the initial price of the television is $3,000, what is the final amount Rachel will pay?

OpenStudy (mathmale):

Alex: What does "15% off" mean to you? Have you ever gone into a store and purchases something at 20% off? How would you calculate the sale price?

OpenStudy (anonymous):

Divided and I got 15000, and then added 50 and got 15050 but that's not on the answer choices?

OpenStudy (mathmale):

Alex, it'd help me and it'd help you if you's please answer my questions; they were intended to help you get on track.

OpenStudy (anonymous):

Well don't you dividde 3000 and .20.. I said that.

OpenStudy (mathstudent55):

To find a percent of a number, _multiply_ that percent by the number. First, find 15% of $3000.

OpenStudy (mathstudent55):

@alexDelicruze Are you still there?

OpenStudy (anonymous):

Yes. 450

OpenStudy (mathstudent55):

Correct. Since it's a discount, subtract 450 from 3000. What do you get?

OpenStudy (anonymous):

Initial price is \(p_i\). Final price is \(p_f\). \[ p_f = p_i\times (1-15\%)\times (1-5\%) - \$50 \]Well it says that \(p_i = \$3,000\) so... \[ p_f = \$3,000\times (1-15\%)\times (1-5\%) - \$50 \]

OpenStudy (anonymous):

2550

OpenStudy (mathstudent55):

Correct. Now there is a 5% discount on the sale price. Find 5% of $2550.

OpenStudy (anonymous):

1275

OpenStudy (mathstudent55):

Be careful. 5% = 0.05, not 0.50

OpenStudy (anonymous):

127.5

OpenStudy (mathstudent55):